OpenStudy (anonymous):

Please help.... what is the difference between slope-intercept form and slope- intercept formula?

OpenStudy (anonymous):

slope intercept is when you know the y intercept slope intercept is y = mx+b where m is the slope and b the y intercept point slope form is when you have 2 given points formula: y = y(1) + m+ ( x - x(1) ) they are basically the same think its just that one is easier to use than the other .... it depends on the problem

OpenStudy (anonymous):

become a fan ? :D

OpenStudy (anonymous):

so if the question says Take the points you have selected and write the equation of the line using the slope-intercept formula. Show the work. and my points are(2, 4) and (1,-2). oh and I became a fan

OpenStudy (anonymous):

how would I do this

OpenStudy (anonymous):

thank you , and one second ill show you

OpenStudy (anonymous):

The equation is Y = Mx + B. In this equation, m is the slope value and B is the y intercept. X and Y are coordinates of the line which can be substituted into the equation to solve for other variables. Example of equation: Y = 3x + 7.

OpenStudy (anonymous):

oh and my slope is 6

OpenStudy (anonymous):

yes m=slope

OpenStudy (anonymous):

so would it be y = 6x - 4 ? and which point am I supposed to use

OpenStudy (ranga):

Point Slope form: (y - y1) = m(x - x1) m = 6 For the point (x1, y1) you can use either of the two points: (2,4) or (1,-2). After you simplify and express the answer in the Slope-Intercept form, y = mx + b, you will see you get identical results no matter which point you use.

OpenStudy (ranga):

Do you want me to complete this problem?

OpenStudy (anonymous):

yes please

OpenStudy (ranga):

Okay. I will start from the beginning. You are given two points on the line: (2,4) and (1,-2) Given two points you can quickly find the slope. Slope m = (y1-y2)(x1-x2) = (4-(-2)) / (2-1) = 6 They want the answer in the slope-intercept formula. That is, y = mx + b Put the value for m. y = 6x + b We need to solve for b. You can put either one of the points into the equation. Let us take the first point (2,4). Put x = 2 and y = 4 into the equation and solve for b: y = 6x + b 4 = (6)(2) + b 4 = 12 + b b = 4 - 12 = -8 y = 6x - 8 is the equation of the line in slope-intercept form.

OpenStudy (ranga):

We could also have chosen the second point and we will still end up with the same answer. Let us start with y = 6x + b and try the second point (1,-2) put x = 1 and y = -2 in y = 6x + b -2 = 6(1) + b -2 = 6 + b -2 - 6 = b b = -8 The equation is y = 6x - 8 (same as before).

OpenStudy (ranga):

The -8 is called the y-intercept. It is where the line crosses the y-axis. The graph of the line will look like this: